Suppose a countrys population grows by 2% a year and at the same time its real GDP shrinks by 1% a year. Real GDP per person is increasing by- a year. a. -3% b. 2% c. 1% d.190 3%
Answer #1

Real GDP per person is given by dividing total real GDP by the number of persons comprising country's population.

RGDPP RGDP

Let initial RGDPi be R0 and initial population be N0.

Intial RGDP per person

RGDP Ro No RGDPP

After a year, New RGDP = RGDPf = R0 + %change in R0

= R0 - 1% of R0 = R0 - 0.01R0 = 0.99R0

New population = Nf = N0 + %change in N0

= N0 + 2% of N0 = N0 + 0.02N0 = 1.02N0

Thus, new real GDP per person

RGDP 0.99R0 0.99Ro 1 .0277) 0.9710 No Ro = RG DPP,- =

Therefore % change in RGDPP =

0.97翯-No RG DPP, _ RG D PP, RGDPP 102(グ

Thus real GDP per person is increasing by -3% per year (i.e. it is actually shrinking by 3% per year).